OpenStudy (anonymous):

Please help me solve the problem 4Y+17=21

OpenStudy (anonymous):

4y = 21-17 y = 4/4 y = 1

OpenStudy (anonymous):

I understand how you got the answer but doesnt that seem advanced for a fourth grader

OpenStudy (anonymous):

no